# Exponential Convergence of Augmented Primal-dual Gradient Algorithms for Partially Strongly Convex Functions

### v1

@@ V-B Relaxing global strong convexity
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})$
$\geq$ $\mu\|x-x_{p}\|^{2}\geq\frac{\mu}{l^{2}}\|\nabla f(x)-\nabla f(x_{p})\|^{2}.$ (24)
@@
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})$
$\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x_{p})\|^{2}$ (25)
@@
Define $\phi(x)=f(x)-\nabla f(x_{0})^{\top}x$, for a fixed $x_{0}\in\mathbb{R}^{n}$. It is clear that $\phi(x)$ has $l$-Lipschitz gradient. By Assumption 7, $\phi(x)$ has a global optimal point $x=x_{0}$. Then, following similar arguments from [30, Theorem 2.1.5], we conclude (2). ∎
@@
This generalization broadens the applicability of exponential convergence results to problems that do not satisfy conventional convexity conditions. Our result coincides with [17] and is equivalent to the EB condition in [15] under convexity assumption [6].
@@ VI Conclusion
In this paper, we established the global exponential convergence of augmented primal-dual gradient algorithms for optimization problems with partially strongly convex functions. We demonstrated that the globally strong convexity of the objective function can be relaxed to strong convexity only in the subspace defined by the equality constraints, provided that the global Lipschitz condition for the gradient is satisfied. This relaxation allows the algorithm to achieve exponential convergence even when the objective function is merely convex outside this subspace. Future work could explore further extensions of these results to optimization problems with inequality constraints.

### v2

@@ V-B Relaxing global strong convexity
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})\geq\mu\|x-x_{p}\|^{2}$
$\geq\frac{\mu}{l^{2}}\|\nabla f(x)-\nabla f(x_{p})\|^{2}.$ (18)
@@
$\left(\nabla f(x)-\nabla f(x_{p})\right)^{\top}(x-x_{p})\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x_{p})\|^{2}$ (19)
@@
Define $\phi(x)=f(x)-\nabla f(x_{0})^{\top}x$, for a fixed $x_{0}\in\mathbb{R}^{n}$. It is clear that $\phi(x)$ has $l$-Lipschitz gradient. By Assumption 7, $\phi(x)$ has a global optimal point $x=x_{0}$. Then, following similar arguments from [32, Theorem 2.1.5], we conclude (19). ∎
@@
This generalization broadens the applicability of exponential convergence results to problems that do not satisfy convex conditions. Our result coincides with [17] and is equivalent to the EB condition in [15] under the convexity assumption [6].
@@ VI Conclusion
This paper established the global exponential convergence of augmented primal-dual gradient algorithms for optimization problems with partially strongly convex functions. The globally strong convexity of the objective function can be relaxed to strong convexity only in the subspace defined by the equality constraints, provided that the global Lipschitz condition for the gradient is satisfied. This relaxation allows the algorithm to achieve exponential convergence even when the objective function is merely convex outside this subspace. Future work could further extend these results to optimization problems with inequality constraints.